OpenStudy (anonymous):

How do I graph these (Algebra II) please? 1) y = x^2 + 4x + 1 2) f(x) = 3x^2 - 4x - 2

jimthompson5910 (jim_thompson5910):

You can plot a series of points and draw a curve through them or you can plot the vertex and and a point on the parabola

OpenStudy (anonymous):

but how do i do that?

jimthompson5910 (jim_thompson5910):

to get any point, you plug in any number for x to get a corresponding y value when x = 0, y is... y = x^2 + 4x + 1 y = (0)^2 + 4(0) + 1 ... plug in x = 0 y = 0 + 4(0) + 1 y = 0 + 0 + 1 y = 1 So when x = 0, y is y = 1 giving us the point (0, 1)

OpenStudy (anonymous):

I can do that for both of them?

jimthompson5910 (jim_thompson5910):

When x = 1, y is... y = x^2 + 4x + 1 y = (1)^2 + 4(1) + 1 ... plug in x = 1 y = 1 + 4(1) + 1 y = 1 + 4 + 1 y = 6 So when x = 1, y is y = 6 giving us the point (1, 6)

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you can do this for any value of x to get any point once you have enough points, plot them and draw a curve through them

jimthompson5910 (jim_thompson5910):

do each equation one at a time though

OpenStudy (anonymous):

thank you so much :)

jimthompson5910 (jim_thompson5910):

np